Grade 12th passPhysical Chemistry

A face centred cubic lattice is made up of a single type of atom and two of its corner atoms are left unoccupied per unit cell.The packing fraction of the solid is

Meenakshi
7 Years agoGrade 12th pass

1 Answer

Pooja
7 Years ago
Dear Student,
For this type of cubic lattice Zeff = 3+3/4= 15/4 (no. of atoms in a cubic lattice)
Packing fraction = Volume occupied by the atoms in the cubic lattice/ volume of the Cubic Lattice
For fcc apply edge lenth to radius conversion
and calculate the PF accordingly